can someone help me solve by completing the square? i don't understand
2x^2+10x+11=0

Answer by stanbon(75887)

2x^2+10x+11=0
2(x^2+5x) = -11
x^2+5x+(5/2)^2= -11/2 + 25/4
(x+(5/2))^2 = 3/4
(x+(5/2)) = sqrt3/2 or (x+(5/2)) = -sqrt3/2
x=(-5+sqrt3)/2 or x=(-5-sqrt3)/2
=============
Cheers,
Stan H.